# Learning Adaptive and View-Invariant Vision Transformer with Multi-Teacher Knowledge Distillation for Real-Time UAV Tracking

### v1

@@ 4.2 Comparison with Lightweight Trackers
2) Performance on WebUAV-3M (Zhang et al., 2022a): As illustrated in Table 2, we have conducted a comparison of our AVTrack-DeiT and AVTrack-MD-DeiT with ten lightweight SOTA trackers on WebUAV-3M. Our AVTrack-DeiT surpasses all other lightweight trackers in success rate, achieving a 1.1% improvement over the second-place Aba-ViTrack while maintaining a comparable precision with only a 0.4% difference. Notably, our AVTrack-MD-DeiT experiences less than a 1.0% drop in both precision and success rate, while improving AGX speed by 9.2%. These results further underscore the effectiveness of our approach.
3) Efficiency Comparison: To further demonstrate that our tracker achieves a better trade-off between accuracy and efficiency, we compare our methods against ten SOTA lightweight trackers, in terms of floating point operations (FLOPs), the number of parameters (Params.), and the speed (AGX.FPS) on the Nvidia Jetson AGX Xavier edge device. The results are shown in Table 2. Notably, since our AVTrack-DeiT tracker features adaptive architectures, the FLOPs and parameters range from minimum to maximum values. As observed, both the minimum FLOPS and Params of our AVTrack-DeiT are notably lower than those of all the SOTA trackers, and even its maximum values are lower than those of most SOTA trackers. Furthermore, our AVTrack-MD-DeiT achieves the lowest FLOPs (1.5G) and parameters (5.3M), except for AVTrack-DeiT, while also delivering the fastest speed at 46.1 AGX.FPS. This comparison in terms of computational complexity also underscores the efficiency of our methods.
@@
Figure 5: The qualitative evaluations were performed on five video sequences from the following datasets: DTB70 (Li and Yeung, 2017), UAVDT (Du et al., 2018), VisDrone2018 (Zhu et al., 2018), UAV123@10fps (Mueller et al., 2016), and UAV123 (Mueller et al., 2016). The evaluated sequences include Surfing12, S1607, uav0000180_00050_s, person10, and truck1.
@@
Table 3: Precision (Prec.), success (Succ.), and average GPU speed (Avg.FPS) comparison between our AVTrack-Deit and DL-based tracker on five UAV tracking benchmarks.
@@
Tracker Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Avg.FPS
DiMP 79.2 61.3 78.3 57.4 83.5 63.0 83.1 65.2 85.1 64.7 81.8 62.3 51.9
PrDiMP 84.0 64.3 75.8 55.9 79.8 60.2 87.2 66.5 83.9 64.7 82.1 62.3 53.6
TrSiam 82.7 63.9 88.9 65.0 84.0 63.5 83.9 66.3 85.3 64.9 84.9 64.7 38.3
TranT 83.6 65.8 82.6 63.2 85.9 65.2 85.0 67.1 84.8 66.5 84.4 65.6 53.2
AutoMatch 82.5 63.4 82.1 62.9 78.1 59.6 83.8 64.4 78.1 59.4 80.9 61.9 63.9
SparseTT 82.3 65.8 82.8 65.4 81.4 62.1 85.4 68.8 82.2 64.9 82.8 65.4 32.4
CSWinTT 80.3 62.3 67.3 54.0 75.2 58.0 87.6 70.5 87.1 68.1 79.5 62.6 12.3
SimTrack 83.2 64.6 76.5 57.2 80.0 60.9 88.2 69.2 87.5 69.0 83.1 64.2 72.8
OSTrack 82.7 65.0 85.1 63.4 84.2 64.8 84.7 67.4 83.1 66.1 83.9 65.3 68.4
ZoomTrack 82.0 63.2 77.1 57.9 81.4 63.6 88.4 69.6 88.8 70.0 83.5 64.8 62.7
SeqTrack 85.6 65.5 78.7 58.8 83.3 64.1 86.8 68.6 85.7 68.1 84.0 65.0 32.3
MAT 83.2 64.5 72.9 54.8 81.6 62.2 86.7 68.3 86.9 68.5 82.3 63.6 71.2
ROMTrack 87.2 67.4 81.9 61.6 86.4 66.7 87.4 69.2 85.0 67.8 85.5 66.5 52.3
DCPT 84.0 64.8 76.8 56.9 83.1 64.2 85.7 68.1 86.9 69.1 83.3 64.6 39.2
HIPTrack 88.4 68.6 79.6 60.9 86.7 67.1 89.2 70.5 89.3 70.6 86.6 67.5 32.1
EVPTrack 85.8 66.5 80.6 61.2 84.5 65.8 88.9 70.2 88.7 70.4 85.7 66.8 26.1
AVTrack-DeiT
84.3 65.0 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.1 64.4 256.8
AVTrack-MD-DeiT
84.0 65.2 83.1 60.3 84.9 64.2 82.6 65.2 83.3 65.5 83.6 64.1 310.6
4.3 Comparison with Deep Trackers
@@
To further validate the superiority of our trackers, we compare AVTrack-DeiT and AVTrack-MD-DeiT with 16 deep SOTA trackers, including DiMP (Bhat et al., 2019), PrDiMP (Danelljan et al., 2020), TrSiam (Wang et al., 2021), TransT (Chen et al., 2021a), AutoMatch (Zhang et al., 2021), SparseTT (Fu et al., 2022), CSWinTT (Song et al., 2022), SimTrack (Chen et al., 2022), OSTrack (Ye et al., 2022), ZoomTrack (Kou et al., 2023), SeqTrack (Chen et al., 2023), MAT (Zhao et al., 2023), ROMTrack (Cai et al., 2023), DCPT (Zhu et al., 2024), HIPTrack (Cai et al., 2024), and EVPTrack (Shi et al., 2024) ,across five UAV tracking datasets. The evaluation results are shown in Table 3, which shows precision (Prec.), success rate (Succ.), their average (Avg.), and average GPU speed (Avg.FPS). As shown, our AVTrack-MD-DeiT and AVTrack-DeiT excel by achieving the fastest and second-fastest speeds while maintaining high performance comparable to the SOTA deep trackers. On average, although AVTrack-MD-DeiT and AVTrack-DeiT exhibit slight gaps in average precision and success rate compared to HIPTrack (Cai et al., 2024), EVPTrack (Shi et al., 2024), and ROMTrack (Cai et al., 2023), they are noticeably slower than our methods in terms of GPU speed. Specifically, our AVTrack-MD-DeiT is 9.7, 11.9, and 5.9 times faster than HIPTrack, EVPTrack, and ROMTrack, respectively. These results indicate that our method delivers both high precision and speed, validating its suitability for real-time UAV tracking that prioritizes efficiency as well as accuracy.
@@
Table 4: Comparison of original teacher models and distillation models with varying teacher combinations in terms of performance. $\psi_{1}$, $\psi_{2}$, and $\psi_{3}$ represent the AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA, respectively.
Model $\psi_{1}$ $\psi_{2}$ $\psi_{3}$ DTB70 UAVDT VisDrone UAV123 UAV123@10fps Avg.
@@
AVTrack-DeiT - - - 84.3 65.0 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.1 64.4
AVTrack-ViT - - - 81.3 63.3 79.9 57.7 86.4 65.9 84.0 66.2 83.2 65.7 82.9 63.1
AVTrack-EVA - - - 82.6 64.0 78.8 57.2 84.4 63.5 83.0 64.7 81.2 63.5 82.0 62.6
AVTrack-MD-DeiT $\checkmark$ - - 81.5 62.9 80.5 57.6 82.5 62.4 81.8 64.1 81.1 63.7 81.7↓2.6% 62.1↓2.3%
$\checkmark$ $\checkmark$ - 83.4 64.8 80.9 58.1 83.4 62.9 82.5 64.9 82.4 64.8 82.5↓1.6% 63.1↓1.3%
$\checkmark$ $\checkmark$ $\checkmark$
84.0
65.2
83.1
60.3
84.9
64.2
82.6
65.2
83.3
65.5
83.6
↓0.5%
64.1
↓0.3%
AVTrack-MD-ViT $\checkmark$ - - 79.1 61.2 77.9 55.8 82.1 62.2 81.8 63.8 81.0 62.5 80.3↓2.6% 61.1↓2.0%
$\checkmark$ $\checkmark$ - 81.5 62.9 80.4 58.4 81.8 61.7 82.1 64.2
83.9
65.9
81.9↓1.0% 62.6↓0.5%
$\checkmark$ $\checkmark$ $\checkmark$
84.9
65.7
81.4
59.5
84.8
63.7
82.3
65.1
83.5
65.9
83.4
↑0.6%
64.0
↑0.9%
AVTrack-MD-EVA $\checkmark$ - - 80.1 61.5 77.3 55.8 80.5 61.2 80.6 62.1 80.1 61.9 79.7↓2.3% 60.5↓2.1%
$\checkmark$ $\checkmark$ - 81.2 62.1 79.1 57.1 81.8 62.1
82.1
63.1
80.4 62.1 80.9↓1.1% 61.3↓1.3%
$\checkmark$ $\checkmark$ $\checkmark$
83.2
63.9
80.8
58.0
84.0
63.5
81.5 62.3
82.7
64.7
82.4
↑0.4%
62.5
↓0.1%
Table 5: Comparison of original teacher models and distillation models in terms of model complexity and average GPU speed.
Params. FLOPs Avg.FPS
AVTrack-DeiT 3.5-8.0 0.97-2.4 250.2
AVTrack-ViT 2.4-5.8 0.67-1.7 283.7
AVTrack-EVA 3.5-7.9 0.97-2.4 256.8
AVTrack-MD-DeiT 5.3 1.5
310.6
↑20.9%
AVTrack-MD-ViT 5.3 1.5
303.1
↑21.1%
AVTrack-MD-EVA 3.7 1.1
334.4
↑17.9%
4.4 Ablation Study
Impact of the Number of Teachers: More teachers may provide more diverse knowledge. Table 4 summarizes the performance of the teacher models and distillation models with various teacher combinations, while Table 5 presents their model complexity and average GPU speed. As shown, single-teacher distillation can significantly reduce model complexity during inference and achieve a notable speedup of 17%. However, it leads to a considerable decline in performance compared to the corresponding teacher model, with decreases of over 2.0% in average precision and success rate. In contrast, multi-teacher distillation methods, including double-teacher and triple-teacher distillation, not only reduce model complexity and provide a significant speedup but also result in a smaller performance decrease, with all drops remaining below 1.6% in average precision and success rate, and some even surpassing the performance of the teacher models. In the triple-teacher distillation, the student models (i.e., AVTrack-MD-ViT and AVTrack-MD-EVA) exhibit superior performance when compared to their corresponding teacher models with the same backbone. Specifically, AVTrack-MD-ViT achieves gains of 0.5% in average precision and 0.9% in success rate, while AVTrack-MD-EVA shows a 0.4% improvement in average precision, with only a 0.1% drop in success rate. Although AVTrack-MD-DeiT does not outperform AVTrack-DeiT, it experiences only a slight performance drop, with both average precision and success rate decreasing by less than 0.5%. Notably, in practice, the model complexity of AVTrack is higher than that of AVTrack-MD, as the activated blocks in AVTrack typically exceed half of the total blocks in the backbone. As a result, AVTrack is slower than AVTrack-MD when both use the same ViT backbone. These experimental results provide evidence for the effectiveness of our proposed multi-teacher knowledge distillation based on MI Maximization, with the performance gain attributed to the complementarity of multiple teacher models.
Table 6: Performance of varying numbers of ViT blocks in student models across five UAV tracking datasets.
@@ 4.4 Ablation Study
Table 7: Ablation study on the weighting of loss $\mathcal{L}_{MI}$ in training AVTrack-MD-DeiT by varying $\eta$ from $0.1\times 10^{-1}$ to $0.1\times 10^{-6}$.
@@
Impact of varying numbers of ViT blocks in student models: To closely investigate the impact of varying numbers of ViT blocks on performance in student models, we train the AVTrack-MD-DeiT student model using a range of block counts from 4 to 8. The evaluation results are shown in Table 6. As observed, the number of ViT blocks in the student model directly affects both tracking performance and speed. On average, increasing the number of ViT blocks leads to an upward trend in accuracy and model complexity but a decline in tracking speed. As the number of blocks increases from 4 to 6, each additional block results in a greater than 1.0% improvement in both average precision and average success rate, an increase of more than 0.4M in parameters, and a rise of over 0.1G in FLOPs, accompanied by an approximately 10% decrease in speed. When the number of blocks exceeds 6, further increases do not lead to significant performance gains, while model complexity continues to rise substantially and speed continues to decrease significantly. Considering the balance between performance and speed, we have chosen to set the default number of ViT blocks in the student model to 6 in our implementation, which is half that of the teacher models.
Impact of Weighting the MI maximization-based multi-teacher knowledge distillation: To obtain the optimal weight $\eta$ for the proposed MI maximization-based multi-teacher knowledge distillation, we trained AVTrack-MD-DeiT using varied values of $\eta$ ranging from $0.1\times 10^{-1}$ to $0.1\times 10^{-6}$ with with a scale factor of 0.1. Evaluation results are shown in Table 7. As shown, our tracker achieves best performance when the loss weight ($\eta$) is set to $0.1\times 10^{-3}$. Additionally, we have observed that the second- and third-best performances across these datasets are scattered both above and below the value of $0.1\times 10^{-3}$, without any apparent patterns. This variation may be attributed to the inherent differences between these datasets. Setting a value of $0.1\times 10^{-1}$ for the loss weight ($\eta$) results in an average maximal difference of Avg. Prec. of 1.8% and a maximal difference of Avg. Succ. of 1.3%. These significant differences clearly demonstrate that the choice of weight a has a considerable impact on the tracking performance. To be more specific, when the proposed loss is appropriately weighted, it can enhance the tracking performance. However, if not properly weighted, it may have detrimental effects on the training of the tracking task.
Table 8: Comparison of different multi-teache knowledge distillation loss functions shows that the proposed MI maximization-based distillation framework outperforms the MSE loss.
@@
Impact of Multi-Teacher Knowledge Distillation Based on MI Maximization: To demonstrate the superiority of the proposed multi-teacher knowledge distillation based on the MI maximization, we train the model using MSE loss and the proposed MI-based loss for multi-teacher knowledge distillation, respectively. The evaluation results across five datasets are presented in Table 8. As observed, using the MSE loss results in greater performance loss compared to our method. On average, employing the proposed $L_{MI}$ results in a notable enhancement over the MSE loss, with improvements of 1.9% in precision and 1.7% in success rate, respectively. Additionally, compared to the teacher model, our method incurs only a slight decrease of 0.5% in average precision and 0.3% in average success rate, while achieving a notable speedup of 20.9%. These comparisons show that the benefit of our multi-teacher knowledge distillation based on the MI maximization is that it can provide a more thorough measure of the relationship between features through MI-based loss, which enables the student model to learn a more effective representation from the teacher model. On the other hand, our MI-based loss is less sensitive to noise and outliers compared to MSE, making it particularly effective in noisy environments.
4.5 Qualitative Results:

### v2

@@ IV-B Comparison with Lightweight Trackers
2) Performance on WebUAV-3M [105]: As illustrated in Table II, we have conducted a comparison of our AVTrack-DeiT and AVTrack-MD-DeiT with ten lightweight SOTA trackers on WebUAV-3M. Our AVTrack-DeiT surpasses all other lightweight trackers in success rate, achieving a 1.1% improvement over the second-place Aba-ViTrack while maintaining a comparable precision with only a 0.4% difference. Notably, our AVTrack-MD-DeiT experiences less than a 1.0% drop in both precision and success rate, while improving AGX speed by 9.2%. These results further underscore the effectiveness of our approach.
3) Efficiency Comparison: To further demonstrate that our tracker achieves a better trade-off between accuracy and efficiency, we compare our methods against ten SOTA lightweight trackers, in terms of floating point operations (FLOPs), the number of parameters (Params.), and the speed (AGX.FPS) on the Nvidia Jetson AGX Xavier edge device. The results are shown in Table II. Notably, since our AVTrack-DeiT tracker features adaptive architectures, the FLOPs and parameters range from minimum to maximum values. As observed, both the minimum FLOPS and Params of our AVTrack-DeiT are notably lower than those of all the SOTA trackers, and even its maximum values are lower than those of most SOTA trackers. Furthermore, our AVTrack-MD-DeiT achieves the lowest FLOPs (1.5G) and parameters (5.3M), except for AVTrack-DeiT, while also delivering the fastest speed at 46.1 AGX.FPS. This comparison in terms of computational complexity also underscores the efficiency of our methods.
@@
Figure 5: The qualitative evaluations were performed on five video sequences from the following datasets: DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123@10fps [88], and UAV123 [88]. The evaluated sequences include Surfing12, S1607, uav0000180_00050_s, person10, and truck1.
@@
TABLE III: Precision (Prec.), success (Succ.), and average GPU speed (Avg.FPS) comparison between our AVTrack-Deit and DL-based tracker on five UAV tracking benchmarks.
@@
Tracker Source Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Prec. Succ. Avg.FPS
DiMP[111] ICCV 19 79.2 61.3 78.3 57.4 83.5 63.0 83.1 65.2 85.1 64.7 81.8 62.3 51.9
PrDiMP[112] CVPR 20 84.0 64.3 75.8 55.9 79.8 60.2 87.2 66.5 83.9 64.7 82.1 62.3 53.6
TrSiam[113] CVPR 21 82.7 63.9 88.9 65.0 84.0 63.5 83.9 66.3 85.3 64.9 84.9 64.7 38.3
TranT[114] CVPR 21 83.6 65.8 82.6 63.2 85.9 65.2 85.0 67.1 84.8 66.5 84.4 65.6 53.2
AutoMatch[115] ICCV 21 82.5 63.4 82.1 62.9 78.1 59.6 83.8 64.4 78.1 59.4 80.9 61.9 63.9
SparseTT[116] IJCAI 22 82.3 65.8 82.8 65.4 81.4 62.1 85.4 68.8 82.2 64.9 82.8 65.4 32.4
CSWinTT[117] CVPR 22 80.3 62.3 67.3 54.0 75.2 58.0 87.6 70.5 87.1 68.1 79.5 62.6 12.3
SimTrack[13] ECCV 22 83.2 64.6 76.5 57.2 80.0 60.9 88.2 69.2 87.5 69.0 83.1 64.2 72.8
OSTrack[12] ECCV 22 82.7 65.0 85.1 63.4 84.2 64.8 84.7 67.4 83.1 66.1 83.9 65.3 68.4
ZoomTrack[16] NIPS 23 82.0 63.2 77.1 57.9 81.4 63.6 88.4 69.6 88.8 70.0 83.5 64.8 62.7
SeqTrack[118] CVPR 23 85.6 65.5 78.7 58.8 83.3 64.1 86.8 68.6 85.7 68.1 84.0 65.0 32.3
MAT[119] CVPR 23 83.2 64.5 72.9 54.8 81.6 62.2 86.7 68.3 86.9 68.5 82.3 63.6 71.2
ROMTrack[120] ICCV 23 87.2 67.4 81.9 61.6 86.4 66.7 87.4 69.2 85.0 67.8 85.5 66.5 52.3
DCPT[121] ICRA 24 84.0 64.8 76.8 56.9 83.1 64.2 85.7 68.1 86.9 69.1 83.3 64.6 39.2
HIPTrack[19] CVPR 24 88.4 68.6 79.6 60.9 86.7 67.1 89.2 70.5 89.3 70.6 86.6 67.5 32.1
EVPTrack[18] AAAI 24 85.8 66.5 80.6 61.2 84.5 65.8 88.9 70.2 88.7 70.4 85.7 66.8 26.1
AVTrack-DeiT Ours 84.3 65.0 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.1 64.4 256.8
AVTrack-MD-DeiT Ours 84.0 65.2 83.1 60.3 84.9 64.2 82.6 65.2 83.3 65.5 83.6 64.1 310.6
IV-C Comparison with Deep Trackers
@@
To further validate the superiority of our trackers, we compare AVTrack-DeiT and AVTrack-MD-DeiT with 16 deep SOTA trackers, including DiMP [111], PrDiMP [112], TrSiam [113], TransT [114], AutoMatch [115], SparseTT [116], CSWinTT [117], SimTrack [13], OSTrack [12], ZoomTrack [16], SeqTrack [118], MAT [119], ROMTrack [120], DCPT [121], HIPTrack [19], and EVPTrack [18] ,across five UAV tracking datasets. The evaluation results are shown in Table III, which shows precision (Prec.), success rate (Succ.), their average (Avg.), and average GPU speed (Avg.FPS). As shown, our AVTrack-MD-DeiT and AVTrack-DeiT excel by achieving the fastest and second-fastest speeds while maintaining high performance comparable to the SOTA deep trackers. On average, although AVTrack-MD-DeiT and AVTrack-DeiT exhibit slight gaps in average precision and success rate compared to HIPTrack [19], EVPTrack [18], and ROMTrack [120], they are noticeably slower than our methods in terms of GPU speed. Specifically, our AVTrack-MD-DeiT is 9.7, 11.9, and 5.9 times faster than HIPTrack, EVPTrack, and ROMTrack, respectively. These results indicate that our method delivers both high precision and speed, validating its suitability for real-time UAV tracking that prioritizes efficiency as well as accuracy.
@@
TABLE IV: Comparison of original teacher models and distillation models with varying teacher combinations: model complexity, performance, and average GPU speed. $\psi_{1}$, $\psi_{2}$, and $\psi_{3}$ represent the AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA, respectively.
Model $\psi_{1}$ $\psi_{2}$ $\psi_{3}$ Params FLOPS DTB70 UAVDT VisDrone UAV123 UAV123@10fps Avg. Avg.FPS
@@
AVTrack-DeiT - - - 3.5-7.9 0.97-2.4 84.3 65.0 82.1 58.7 86.0 65.3 84.8 66.8 83.2 65.8 84.1 64.4 256.8
AVTrack-ViT - - - 3.5-8.0 0.97-2.4 81.3 63.3 79.9 57.7 86.4 65.9 84.0 66.2 83.2 65.7 82.9 63.1 250.2
AVTrack-EVA - - - 2.4-5.8 0.67-1.7 82.6 64.0 78.8 57.2 84.4 63.5 83.0 64.7 81.2 63.5 82.0 62.6 283.7
AVTrack-MD-DeiT $\checkmark$ - - 5.3 1.5 81.5 62.9 80.5 57.6 82.5 62.4 81.8 64.1 81.1 63.7 81.7↓2.6% 62.1↓2.3%
$\checkmark$ $\checkmark$ - 83.4 64.8 80.9 58.1 83.4 62.9 82.5 64.9 82.4 64.8 82.5↓1.6% 63.1↓1.3% 310.6↑20.9%
$\checkmark$ $\checkmark$ $\checkmark$ 84.0 65.2 83.1 60.3 84.9 64.2 82.6 65.2 83.3 65.5 83.6↓0.5% 64.1↓0.3%
AVTrack-MD-ViT $\checkmark$ - - 5.3 1.5 79.1 61.2 77.9 55.8 82.1 62.2 81.8 63.8 81.0 62.5 80.3↓2.6% 61.1↓2.0%
$\checkmark$ $\checkmark$ - 81.5 62.9 80.4 58.4 81.8 61.7 82.1 64.2 83.9 65.9 81.9↓1.0% 62.6↓0.5% 303.1↑21.1%
$\checkmark$ $\checkmark$ $\checkmark$ 84.9 65.7 81.4 59.5 84.8 63.7 82.3 65.1 83.5 65.9 83.4↑0.6% 64.0↑0.9%
AVTrack-MD-EVA $\checkmark$ - - 3.7 1.1 80.1 61.5 77.3 55.8 80.5 61.2 80.6 62.1 80.1 61.9 79.7↓2.3% 60.5↓2.1%
$\checkmark$ $\checkmark$ - 81.2 62.1 79.1 57.1 81.8 62.1 82.1 63.1 80.4 62.1 80.9↓1.1% 61.3↓1.3% 334.4↑17.9%
$\checkmark$ $\checkmark$ $\checkmark$ 83.2 63.9 80.8 58.0 84.0 63.5 81.5 62.3 82.7 64.7 82.4↑0.4% 62.5↓0.1%
IV-D Ablation Study
Impact of the Number of Teachers: More teachers may provide more diverse knowledge. Table IV summarizes the model complexity, performance, and average GPU speed of the original teacher models and the distillation models using different teacher combinations. As shown, single-teacher distillation can significantly reduce model complexity during inference and achieve a notable speedup of 17%. However, it leads to a considerable decline in performance compared to the corresponding teacher model, with decreases of over 2.0% in average precision and success rate. In contrast, multi-teacher distillation methods, including double-teacher and triple-teacher distillation, not only reduce model complexity and provide a significant speedup but also result in a smaller performance decrease, with all drops remaining below 1.6% in average precision and success rate, and some even surpassing the performance of the teacher models. In the triple-teacher distillation, the student models (i.e., AVTrack-MD-ViT and AVTrack-MD-EVA) exhibit superior performance when compared to their corresponding teacher models with the same backbone. Specifically, AVTrack-MD-ViT achieves gains of 0.5% in average precision and 0.9% in success rate, while AVTrack-MD-EVA shows a 0.4% improvement in average precision, with only a 0.1% drop in success rate. Although AVTrack-MD-DeiT does not outperform AVTrack-DeiT, it experiences only a slight performance drop, with both average precision and success rate decreasing by less than 0.5%. Notably, in practice, the model complexity of AVTrack is higher than that of AVTrack-MD, as the activated blocks in AVTrack typically exceed half of the total blocks in the backbone. As a result, AVTrack is slower than AVTrack-MD when both use the same ViT backbone. These experimental results provide evidence for the effectiveness of our proposed multi-teacher knowledge distillation based on MI Maximization, with the performance gain attributed to the complementarity of multiple teacher models.
TABLE V: Performance of varying numbers of ViT blocks in student models across five UAV tracking datasets.
@@ IV-D Ablation Study
TABLE VI: Ablation study on the weighting of loss $\mathcal{L}_{MI}$ in training AVTrack-MD-DeiT by varying $\eta$ from $0.1\times 10^{-1}$ to $0.1\times 10^{-6}$.
@@
Impact of varying numbers of ViT blocks in student models: To closely investigate the impact of varying numbers of ViT blocks on performance in student models, we train the AVTrack-MD-DeiT student model using a range of block counts from 4 to 8. The evaluation results are shown in Table V. As observed, the number of ViT blocks in the student model directly affects both tracking performance and speed. On average, increasing the number of ViT blocks leads to an upward trend in accuracy and model complexity but a decline in tracking speed. As the number of blocks increases from 4 to 6, each additional block results in a greater than 1.0% improvement in both average precision and average success rate, an increase of more than 0.4M in parameters, and a rise of over 0.1G in FLOPs, accompanied by an approximately 10% decrease in speed. When the number of blocks exceeds 6, further increases do not lead to significant performance gains, while model complexity continues to rise substantially and speed continues to decrease significantly. Considering the balance between performance and speed, we have chosen to set the default number of ViT blocks in the student model to 6 in our implementation, which is half that of the teacher models.
Impact of Weighting the MI maximization-based multi-teacher knowledge distillation: To obtain the optimal weight $\eta$ for the proposed MI maximization-based multi-teacher knowledge distillation, we trained AVTrack-MD-DeiT using varied values of $\eta$ ranging from $0.1\times 10^{-1}$ to $0.1\times 10^{-6}$ with with a scale factor of 0.1. Evaluation results are shown in Table VI. As shown, our tracker achieves best performance when the loss weight ($\eta$) is set to $0.1\times 10^{-3}$. Additionally, we have observed that the second- and third-best performances across these datasets are scattered both above and below the value of $0.1\times 10^{-3}$, without any apparent patterns. This variation may be attributed to the inherent differences between these datasets. Setting a value of $0.1\times 10^{-1}$ for the loss weight ($\eta$) results in an average maximal difference of Avg. Prec. of 1.8% and a maximal difference of Avg. Succ. of 1.3%. These significant differences clearly demonstrate that the choice of weight a has a considerable impact on the tracking performance. To be more specific, when the proposed loss is appropriately weighted, it can enhance the tracking performance. However, if not properly weighted, it may have detrimental effects on the training of the tracking task.
TABLE VII: Comparison of different multi-teache knowledge distillation loss functions shows that the proposed MI maximization-based distillation framework outperforms the MSE loss.
@@
Impact of Multi-Teacher Knowledge Distillation Based on MI Maximization: To demonstrate the superiority of the proposed multi-teacher knowledge distillation based on the MI maximization, we train the model using MSE loss and the proposed MI-based loss for multi-teacher knowledge distillation, respectively. The evaluation results across five datasets are presented in Table VII. As observed, using the MSE loss results in greater performance loss compared to our method. On average, employing the proposed $L_{MI}$ results in a notable enhancement over the MSE loss, with improvements of 1.9% in precision and 1.7% in success rate, respectively. Additionally, compared to the teacher model, our method incurs only a slight decrease of 0.5% in average precision and 0.3% in average success rate, while achieving a notable speedup of 20.9%. These comparisons show that the benefit of our multi-teacher knowledge distillation based on the MI maximization is that it can provide a more thorough measure of the relationship between features through MI-based loss, which enables the student model to learn a more effective representation from the teacher model. On the other hand, our MI-based loss is less sensitive to noise and outliers compared to MSE, making it particularly effective in noisy environments.
IV-E Qualitative Results: